# Scrambled Halton Subsequences and Inverse Star-Discrepancy

### v1

@@ 3 Theoretical background
Let $(x_{n})$ be a sequence of points in $[0,1)^{d-1}$ with discrepancy $D_{N}^{*}(x_{n})$. Then it is possible to construct a point set in $[0,1)^{d}$ which has almost the same $d$-dimensional discrepancy. By defining $y_{n}=(n/N,x_{n})\in[0,1)^{d}$ for $n=1,\ldots,N-1$, we obtain
$D_{N}^{*}(P)\leq\max_{1\leq M\leq N}\frac{M}{N}D_{M}^{*}(x_{n})+\frac{1}{N}\leq c_{d-1}\frac{\log(N)}{N}+\frac{1}{N},$ (3)
see [Nie92], Lemma 3.7. For the Halton sequence, this specific point set is known as Hammersley point set. Vice versa, if a bound of the form $D_{N}^{*}(x_{n})\leq c_{d}\log(N)^{d-1}/N$ holds for all point sets in dimension $d$, then a bound of the form $D_{N}^{*}(x_{n})\leq c_{d-1}^{\prime}\log(N)^{d-1}/N$ is true for all sequences in $[0,1)^{d-1}$, see [Rot54].
@@
The concept of bracketing numbers serves as an important tool for finding bounds on the discrepancy, compare e.g. [GPW21] and [DGS05]. Here we follow the definition given in [Gne08]: Let $A\subset[0,1]^{d}$ and $\delta\in(0,1]$. A finite set of points $\Gamma\subset[0,1]^{d}$ is called a $\delta$-cover of $A$, if for every $y\in A$, there exist $x,z\in\Gamma\cup\left\{0\right\}$ such that $x\leq y\leq z$ (to be read component-wise) and $\lambda_{d}([0,z])-\lambda_{d}([0,x])\leq\delta$, where $\lambda_{d}$ denotes the $d$-dimensional Lebesgue-measure. For $x\leq y$ the interval $[x,y]$ is defined by $[x,y]:=[x_{1},y_{1}]\times[x_{2},y_{2}]\times\ldots\times[x_{d},y_{d}]$ and similarly for half-open intervals. The bracketing number $N_{[\ ]}(A,\delta)$ of $A$ is the smallest number of closed axis-parallel boxes (or brackets) of the form $[x,y]$ with $x,y\in[0,1]^{d}$, satisfying $\lambda_{d}([0,y])-\lambda_{d}([0,x])\leq\delta$, whose union contains $A$. Here, we will use bracketing numbers for proving Theorem 4.
@@
The key ingredient for the proof is the following Theorem 8 which is to a great extent analogous to Theorem 3.5 in [GPW21]. The main difference is the restriction to the case $d\geq 5$. The remaining cases $d\leq 4$ in the proof of Theorem 4 can then be covered by Halton sequences. Some theoretical work is also required for these small dimensions in order to speed up the calculations on the computer.
If $N$ is big in comparison to $d\leq 4$, then the theoretical bounds for the Halton sequence/Hammersley point set as a low-discrepancy sequence/point set yields a better bound than the one in Theorem 4 anyhow. Moreover, it is possible to find sets with a sufficiently small star-discrepancy in the finitely many exceptions.
@@
In order to derive Theorem 4 we proceed as in [GPW21] and prove the following result from which it will follow easily for $d\geq 5$.
@@
holds with probability at least $1-e^{-(1.5558085c^{2}-9.4564395)\cdot d}$ implying that for every $q\in(0,1)$
$D_{N}^{*}(X)\leq 0.801718\sqrt{9.456395+\frac{\log\left((1-q)^{-1}\right)}{d}}\cdot\sqrt{\frac{d}{N}}$
@@
Now we can formulate the following lemma.
@@
In our situation the two variables from the lemma are $\sigma=\mu-\log(2^{\mu})-1-\frac{\log(2)}{d}$ and $\zeta=1+\log(2)+\frac{\log(2)}{d}$. We now want to make sure that the expression in the bracket of (4) is $\leq\sqrt{\frac{\pi d}{2}}$. For $\mu=13$ and $\tau_{\mu}=\log(2)/(\mu-\sigma)+0.02120108$ this inequality is satisfied and $c$ is equal to the claimed value. ∎
@@
$D_{N}(\varphi_{2}(n))\leq\frac{7}{2N}+\frac{1}{2\log(2)}\frac{\log(N)}{N}$
@@
$D_{N}(\varphi_{2,3}(n))\leq\frac{3}{N}+\frac{1}{N}\left(\frac{1}{2\log(2)}\log(N)+\frac{3}{2}\right)\left(\frac{1}{\log(3)}\log(N)+2\right)$
which is smaller that $2.463\sqrt{\frac{3}{N}}$ for $N>28$. However for $N\leq 28$, the actual discrepancy of the Hammersley can be calculated with the help of a computer (or also by hand if desired) to see that the claimed inequality is actually true for all $N\in\mathbb{N}$.
For the case $d=4$ the mentioned theoretical bounds on the discrepancy of the Halton sequences only guarantees $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{4}{N}}$ for $N\gg 3\cdot 10^{5}$ which means that it is not feasible to calculate the missing discrepancies for these $N$, even with a very fast computer. However, there are improved bounds on the discrepancy of Halton sequences according to Theorem 2.1 from [Ata04], see also [FKP15].
Theorem 10 (Atanassov).
@@
Applying Theorem 10 and (3) implies that the Hammersley point set satisfies $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{4}{N}}$ for $N>11,759$. The finitely many (relative few) excluded point sets can now be checked on a computer to satisfy the desired inequality. This completes the proof. ∎
@@
Note that our method could be extended to higher dimensions. For the case $d=5$ the mentioned theoretical bounds on the discrepancy of the Halton sequences guarantee $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{5}{N}}$ for $N>7,800$. Again, these finitely many exceptions could be checked on a computer. However, the running time of the DEM algorithm is $O(N^{d/2+1})$, see [CVJ+23]. Therefore, these finitely many exception become less and less feasible to check as the dimension increases.
@@
The following triangle inequality for the discrepancy helps to speed up the calculations of the discrepancy for the many $N$ under consideration: Suppose that $D_{N_{0}}^{*}(X)$ is known and that we want to calculate $D_{N_{1}}^{*}(X)$ for $N_{1}=\alpha N_{0}$ with $\alpha>1$. Then according to [KN74], Theorem 2.6 on p. 115, it holds that
@@
Thus the star-discrepancy can increase by at most $\frac{\alpha-1}{\alpha}$. If $b=2.463\sqrt{\frac{d}{N}}-D_{N_{0}}^{*}(X)$, then we may choose $\alpha=1/(1-b)$. For instance for $N_{0}=5,000$ and $d=4$ we have $D_{N_{0}}^{*}(X)=0.0045$ and $0.06932.463\sqrt{\frac{d}{N}}=$ 0.0693 for the Hammesley point set in base $b=(2,3,5)$. Hence we could directly jump to $N_{1}=5,363$ if we dealt with sequences instead of sets.
Moreover, we know that the star-discrepancy of the underlying Halton sequence is at most $D_{N_{0}}(X)$ as well and the presented argument works for the Halton sequence when moving from $N_{0}$ to $N_{1}$ For the Hammersley point set with $N_{1}$ points, the discrepancy might be bigger by $1/N_{1}<1/N_{0}$ than the one of the Halton sequence according to (3). Hence we may only choose $\alpha=(1-1/N_{0})/(1-b)$. In our numerical example this means that we may only jump only to $,35$ 353. This trick decreases the necessary computational effort by a factor of about $100$.
It was conjectured in [NW21] that $n=10d$ points suffice to reach $d_{\infty}^{*}=0.25$. Our result constitutes partial progress towards this conjecture by a simple application of Theorem 4.
@@
With the value $c=2.4968$ from [GPW21], we would only obtain a set size $N=100d$ for $S_{1}$ and $D_{N}^{*}(S_{2})\leq 0.78956.$.
@@ 4 Appendix
In the appendix, we report the parametrization of the algorithms as well as the scrambled Halton subsequences which yield the best found values for the inverse star-discrepancies (as presented in Tables 2 and 3). In Table 4 and 5 we used the parametrization $n_{\textrm{shifts}}=(100,100,40,40,40)$ and $m_{\pi}=(1,5,20,80,80)$. The data therein needs to be read as follows: at first we list the prime base $b$ for the Halton sequence (these are typically the first $k$ prime numbers). Afterwards we note the applied shifts $x_{i}\mapsto n_{i}\cdot x_{i}$ as vector $n$. Finally, we write down the permutation $\pi_{i}$ for each base element $b_{i}$ in the same ordering as the prime numbers. Thereby $\pi_{i,j}$, the $j$-th entry of $\pi_{i}$, is to be read as $\pi_{i}(j-1)=\pi_{i,j}$.
@@
Regarding the parametrization of the algorithm for obtaining the results in Table 3, we used the same one as for the Tables 4 and 5 in the dimensions $4$ and $5$. Admittedly, finding the optimal parametrization found in Table 6, required a bit more fiddling for dimension $d=7$ because whenever the prime number $p=17$ was used, the discrepancy became larger than in [DGW10]. Besides this peculiarity, the parametrization was $n_{\textrm{shift}}=(100,100,40,40,40,40,40)$ and $m_{\pi}=(1,5,20,80,80,60,60)$. In dimension $d=9$ we used the parametrization $n_{\textrm{shift}}=(100,100,40,40,40,40,40,20,20)$ and $m_{\pi}=(1,5,20,80,80,60,60,60,60)$.
@@
In Figure 1, we plot the star-discrepancy values of the usual van der Corput sequence $\varphi_{3}$, the scrambled van der Corput sequence $\varphi_{3,\pi}$ according to [BW79] and the best found scrambled van der Corput subsequence $\varphi_{3,\pi,1,000}$ for $N=1,000$ each from $n=1$ to $n=1,000$. In Figure 2, the same is done for $p=5$ instead. It becomes clearly visible that the discrepancy values of $\varphi_{\cdot,\pi}$ and $\varphi_{\cdot,\pi,1,000}$ are smaller than the ones of $\varphi_{\cdot}$ for the majority of $n$. The two latter methods achieve the minimal value about the same number of times.

### v2

@@ 3 Theoretical background
Let $(x_{n})$ be a sequence of points in $[0,1)^{d-1}$ with star-discrepancy $D_{N}^{*}(x_{n})$. Then it is possible to construct a point set in $[0,1)^{d}$ which has almost the same $d$-dimensional star-discrepancy. By defining $y_{n}=(n/N,x_{n})\in[0,1)^{d}$ for $n=1,\ldots,N-1$, and $P=\{y_{1},\ldots,y_{N-1}\}$ we obtain
$D_{N}^{*}(P)\leq\max_{1\leq M\leq N}\frac{M}{N}D_{M}^{*}(x_{n})+\frac{1}{N},$
see [Nie92], Lemma 3.7. For the Halton sequence, this specific point set is known as Hammersley point set and the bound becomes
$D_{N}^{*}(P)\leq c_{d-1}\frac{\log(N)^{d-1}}{N}+\frac{1}{N}.$ (3)
Vice versa, if a bound of the form $D_{N}^{*}(P)\leq c_{d}\log(N)^{d-1}/N$ holds for point sets $P$ in dimension $d$, then a bound of the form $D_{N}^{*}(x_{n})\leq c_{d-1}^{\prime}\log(N)^{d-1}/N$ is true for all sequences in $(x_{n})$ in $[0,1)^{d-1}$, see [Rot54].
@@
The concept of bracketing numbers serves as an important tool for finding bounds on the star-discrepancy, compare e.g. [GPW21] and [DGS05]. Here we follow the definition given in [Gne08]: Let $A\subset[0,1]^{d}$ and $\delta\in(0,1]$. A finite set of points $\Gamma\subset[0,1]^{d}$ is called a $\delta$-cover of $A$, if for every $y\in A$, there exist $x,z\in\Gamma\cup\left\{0\right\}$ such that $x\leq y\leq z$ (to be read component-wise) and $\lambda_{d}([0,z])-\lambda_{d}([0,x])\leq\delta$, where $\lambda_{d}$ denotes the $d$-dimensional Lebesgue-measure. For $x\leq y$ the interval $[x,y]$ is defined by $[x,y]:=[x_{1},y_{1}]\times[x_{2},y_{2}]\times\ldots\times[x_{d},y_{d}]$ and similarly for half-open intervals. The bracketing number $N_{[\ ]}(A,\delta)$ of $A$ is the smallest number of closed axis-parallel boxes (or brackets) of the form $[x,y]$ with $x,y\in[0,1]^{d}$, satisfying $\lambda_{d}([0,y])-\lambda_{d}([0,x])\leq\delta$, whose union contains $A$. Here, we will use bracketing numbers for proving Theorem 4.
@@
The main ingredient for the proof is the following Theorem 9 which is to a great extent analogous to Theorem 3.5 in [GPW21]. The main difference is the restriction to the case $d\geq 5$. The remaining cases $d\leq 4$ in the proof of Theorem 4 can then be covered by Halton sequences and some additional ideas to reduce the required number of computer calculations.
If $N$ is big in comparison to $d\leq 4$, then the theoretical bounds for the Halton sequence/Hammersley point set as a low-discrepancy sequence/point set yields a better bound than the one in Theorem 4 anyhow. Moreover, it is possible to find sets with a sufficiently small star-discrepancy in the finitely many exceptions. In small dimensions, we will explain and use a theoretically justifiable trick which helps to speed up the necessary computer calculations, see Remark 13.
@@
In order to derive Theorem 4 we proceed as in [GPW21] and prove the following result from which it follows easily for $d\geq 5$.
@@
holds with probability at least $1-e^{-(1.6728349c^{2}-10.1495427)\cdot d}$ implying that for every $q\in(0,1)$
$D_{N}^{*}(X)\leq 0.7731673\sqrt{10.1495427+\frac{\log\left((1-q)^{-1}\right)}{d}}\cdot\sqrt{\frac{d}{N}}$
@@
Now we use the following lemma, which is essentially a combination of Lemma 3.4 and the beginning of the proof of Theorem 3.5. in [GPW21].
@@
In our situation the two variables from the lemma are $\sigma=\mu-\log(2^{\mu})-1-\frac{\log(2)}{d}$ and $\zeta=1+\log(2)+\frac{\log(2)}{d}$. In order to be able to ignore the exponential term in front of the bracket in (4), we set $c_{0}=\sqrt{(\mu-\sigma)/2}$. We now want to make sure that the expression in the bracket of (4) is $\leq\sqrt{\frac{\pi d}{2}}$. Note that $(\mu-\sigma)$ may be replaced by $2c_{0}^{2}$, which simplifies the calculation. For $\mu=13$ and $\tau_{\mu}=\log(2)/(\mu-\sigma)+0.02120108$ the desired inequality is satisfied and $c$ is equal to the claimed value. ∎
@@
$D_{N}^{*}(\varphi_{2}(n))\leq D_{N}(\varphi_{2}(n))\leq\frac{7}{2N}+\frac{1}{2\log(2)}\frac{\log(N)}{N}$
@@
$D_{N}^{*}(\varphi_{2,3}(n))\leq\frac{3}{N}+\frac{1}{N}\left(\frac{1}{2\log(2)}\log(N)+\frac{3}{2}\right)\left(\frac{1}{\log(3)}\log(N)+2\right)$
which is smaller that $2.463\sqrt{\frac{3}{N}}$ for $N>28$. However, for $N\leq 28$, the actual star-discrepancy of the Hammersley point set can be calculated with the help of a computer (or also by hand if desired) to see that the claimed inequality is actually true for all $N\in\mathbb{N}$.
For the case $d=4$ the mentioned theoretical bounds on the star-discrepancy of the Halton sequences in Theorem 2 only guarantees $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{4}{N}}$ for $N>3\cdot 10^{5}$, even when ignoring the linear term. This means that it is not feasible to calculate the exact star-discrepancies for the missing $N$, even with a very fast computer. However, there are improved bounds on the star-discrepancy of Halton sequences according to Theorem 2.1 from [Ata04], see also [FKP15].
Theorem 11 (Atanassov, [Ata04]).
@@
Applying Theorem 11 and (3) implies that the Hammersley point set satisfies $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{4}{N}}$ for $N>11,759$. The finitely many (relatively few) excluded point sets can now be checked on a computer to satisfy the desired inequality. This completes the proof. ∎
@@
Note that our method could be extended to higher dimensions. For the case $d=5$ the theoretical bounds on the star-discrepancy of the Halton sequences from Theorem 11 guarantee $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{5}{N}}$ for $N>7,800$. Again, these finitely many exceptions could be checked on a computer. However, the running time of the DEM algorithm is $O(N^{d/2+1})$, see [CVN+23]. Therefore, these finitely many exception become less and less feasible to check as the dimension increases.
@@
The following triangle inequality for the star-discrepancy accelerates the computer computation by eliminating the need to calculate star-discrepancies for many values of $N$.: Suppose that $D_{N_{0}}^{*}(X)$ is known and that we want to calculate $D_{N_{1}}^{*}(X)$ for $N_{1}=\alpha N_{0}$ with $\alpha>1$. Then according to [KN74], Theorem 2.6 on p. 115, it holds that
@@
Thus the star-discrepancy can increase by at most $\frac{\alpha-1}{\alpha}$. If $b=2.463\sqrt{\frac{d}{N}}-D_{N_{0}}^{*}(X)$, then we may choose $\alpha=1/(1-b)$. For instance for $N_{0}=5,000$ and $d=4$ we have $D_{N_{0}}^{*}(X)=0.0045$ and $0.06932.463\sqrt{\frac{d}{N}}=$ 0.0693 for the Hammersley point set in base $b=(2,3,5)$. Hence we could directly jump to $N_{1}=5,363$ elements and calculate their discrepancy if we dealt with a sequence. The gain is that it is not necessary to calculate any discrepancy in the range $3625,001\leq N\leq 5,$ 362. This trick decreases the necessary computational effort by a three-digit factor.
For the Hammersley point set, we need however to account for the fact that it is not a sequence: we know that the star-discrepancy of the underlying Halton sequence is at most $D_{N_{0}}^{*}(X)$ as well and the presented argument works for the Halton sequence when moving from $N_{0}$ to $N_{1}$ For the Hammersley point set with $N_{1}$ points, the star-discrepancy might be bigger by $1/N_{1}<1/N_{0}$ than the one of the Halton sequence according to (3). Hence we may only choose $\alpha=(1-1/N_{0})/(1-b)$. In our numerical example this means that we may only jump to $,35$ 353, which is however an acceptable difference.
It was conjectured in [NW21] that $n=10d$ points suffice to reach $D_{N}^{*}=0.25$. Our result constitutes partial progress towards this conjecture by a simple application of Theorem 4.
@@
With the value $c=2.4968$ from [GPW21], we would only obtain a set size $N=100d$ for $S_{1}$ and $D_{N}^{*}(S_{2})\leq 0.78956$.
@@ 4 Appendix
In the appendix, we report the parametrization of the algorithms as well as the scrambled Halton subsequences which yield the best found values for the inverse star-discrepancies (as presented in Tables 2 and 3). In Table 5 and 6 we used the parametrization $n_{\textrm{shifts}}=(100,100,40,40,40)$ and $m_{\pi}=(1,5,20,80,80)$. The data therein needs to be read as follows: at first we list the prime base $b$ for the Halton sequence (these are typically the first $k$ prime numbers). Afterwards we note the applied shifts $x_{i}\mapsto n_{i}\cdot x_{i}$ as vector $n$. Finally, we write down the permutation $\pi_{i}$ for each base element $b_{i}$ in the same ordering as the prime numbers. Thereby $\pi_{i,j}$, the $j$-th entry of $\pi_{i}$, is to be read as $\pi_{i}(j-1)=\pi_{i,j}$.
@@
$N$ Data
$20$ $p=5,n=12,\pi=(0,1,2,3,4)$
$50$ $p=2,n=765,\pi=(0,1)$
$100$ $p=5,n=1,\pi=(0,3,4,1,2)$
$180$ $p=2,n=253,\pi=(0,1)$
$220$ $p=23,n=26$
$\pi=(0,12,9,6,1,10,15,14,7,20,18,3,2,16,5,4,19,13,8,21,17,11,22)$
$260$ $p=2,n=509,\pi=(0,1)$
$350$ $p=2,n=509,\pi=(0,1)$
$420$ $p=2,n=509,\pi=(0,1)$
$500$ $p=2,n=509,\pi=(0,1)$
Table 7: Parameters for the best found scrambled Hammersley point sets in dimension $d=2$
Regarding the parametrization of the algorithm for obtaining the results in Table 3, we used the same one as for the Tables 5 and 6 in the dimensions $4$ and $5$. Admittedly, finding the optimal parametrization found in Table 8, required a bit more fiddling for dimension $d=7$ because whenever the prime number $p=17$ was used, the star-discrepancy became larger than in [DGW10]. Besides this peculiarity, the parametrization was $n_{\textrm{shift}}=(100,100,40,40,40,40,40)$ and $m_{\pi}=(1,5,20,80,80,60,60)$. In dimension $d=9$ we used the parametrization $n_{\textrm{shift}}=(100,100,40,40,40,40,40,20,20)$ and $m_{\pi}=(1,5,20,80,80,60,60,60,60)$. Finally, we present the optimal parametrization for the scrambled Hammersley point sets in dimension $d=2$ in Table 7.
@@
Figure 4 pushes the results from Figure 2 to $N=5,000$ and it becomes visible that $\varphi_{5,\pi,a}$ still has about the same quality as $\varphi_{5,\pi}$ also for $n>1,000$. Both sequences clearly have a smaller discrepancy than the standard van der Corput sequence $\varphi_{5}$.
Figure 4: Comparison of $\varphi_{5},\varphi_{5,\pi}$ and $\varphi_{5,\pi,1,000}$ up to $N=5,000$.